Bob wants to have concrete driveway in front of his house. The driveway will measure 58 2/3 feet long and 18 1/2 feet wide; and will be 3/4 foot thick. How many cubic yards of cement will be needed for the driveway? Round your answer to the nearest cubic yard. [1 cubic yd = 27 cubic feet]
Answer by Alan3354(21605)

Bob wants to have concrete driveway in front of his house. The driveway will measure 58 2/3 feet long and 18 1/2 feet wide; and will be 3/4 foot thick. How many cubic yards of cement will be needed for the driveway? Round your answer to the nearest cubic yard.
-------------------
58 2/3 feet = 704 inches
18.5 feet = 222 inches
3/4 foot = 9 inches
Volume = 704*222*9 cubic inches
V = 1406592 cubic inches
V = 814 cubic feet (/1728)
V = ~ 30 cubic yards (/27)
-------------
English units are stupid and backward.